OpenStudy (anonymous):

Which of the following are trigonometric identities? a. tanx cosx cscx = 1 b. ((secx - cosx)/(secx))= sin^2x c. 1 - tanx tany = (cos(x+y))/ (cosx cosy) d. 4cosx sinx = 2cosx + 1 - 2sinx

OpenStudy (solomonzelman):

tanx cosx cscx = 1 (sinx/cosx)cos(x)(1/sinx)=1 cos and sin cancels and you get 1=1

OpenStudy (anonymous):

Any others? There is supposed to be more than one answer...

OpenStudy (anonymous):

Oh, and thank you for explaining @SolomonZelman.

OpenStudy (solomonzelman):

Anytime

OpenStudy (solomonzelman):

b. ((secx - cosx)/(secx))= sin^2x secx/secx - cosx/secx = sin^2x 1 - cos^2x = sin^2x thus this is also an identity

OpenStudy (anonymous):

None of the other ones? @SolomonZelman

OpenStudy (solomonzelman):

No, I don't think so. You might want to still check them.

OpenStudy (anonymous):

I'm not sure how to. I have a formula sheet here, but I'm not sure how to use it @SolomonZelman

OpenStudy (solomonzelman):

what is tanx ± tany ? I think it might be that one side of choice d is nominator and the other is the denominator or the other way.

OpenStudy (anonymous):

Huh?

OpenStudy (solomonzelman):

I can't work all of them out like that. It's too much.

OpenStudy (anonymous):

I have never learned about \[tanx \pm tany\]...

OpenStudy (solomonzelman):

I mean tanx + tany and tanx - tany

OpenStudy (anonymous):

I know, but I didn't know that tan could have both an x and a y.

OpenStudy (solomonzelman):

We know this cos(x + y) = cosxcosy - sinxsiny => cosxcosy - sinxsiny / cosxcosy = cosxcosy/cosxcosy - sinxsiny/cosxcosy = 1 - tanxtany hence C is an identitity

OpenStudy (solomonzelman):

4cosx sinx = 2cosx + 1 - 2sinx=0 4cos(x)sin(x)-2cos(x)+2sin(x)=0 (2sinx-1)(2cosx+1)=0 2sinx-1=0, sinx=1/2 2cosx+1=0, cosx= -1/2

OpenStudy (solomonzelman):

It's not an identity...

OpenStudy (anonymous):

Ok, thank you so much!

OpenStudy (anonymous):

Are you sure that b is an identity?

OpenStudy (solomonzelman):

yes